A person is 1.7 m height and in a bending position experiences upward acceleration. If the blood pressure difference between the brain and the heart is 8.8 kPa, what is the acceleration that he was experienced? (rblood = 1.1 g/cm3, height difference between heart and brain is 40 cm).

1
Expert's answer
2021-12-27T07:53:32-0500

1) Write the expression that defines how pressure depends on density, acceleration, and height:


"p=\rho hg_\text{total},\\
p=\rho h(g+a)."


2) Express the acceleration "a" and calculate the value of the acceleration:


"p=\rho h(g+a),\\
p=\rho hg+\rho ha,\\\space\\
a=\dfrac{p}{\rho h}-g,\\\space\\
a=\dfrac{8800}{1100 ·0.4}-9.8=10.2\text{ m/s}^2."